OpenStudy (anonymous):

Solve the formula. 7x-8=m;x

OpenStudy (anonymous):

7x-8+8=m+8 7x=m+8 can you continue??

OpenStudy (anonymous):

7x-8=m/x?

OpenStudy (anonymous):

is it 1x?

OpenStudy (anonymous):

@mustry

OpenStudy (anonymous):

yap